# Diffusion-based subsurface multiphysics monitoring and forecasting

### v2

@@ 3.3 Inversion
In realistic monitoring scenarios, besides the historical information of the multiphysics evolution, we might also have some measurements of the seismic response on the surface, such as the seismic recordings, which can be converted to RTM images [Li and Li, 2021]. RTM images play a crucial role in these instances, acting as a strong conditioner that can constrain and guide the generation of CO2 evolution and the corresponding changes in subsurface elastic properties. With our method, we can do an inversion of the subsurface velocity, density, and CO2 saturation, given the observed seismic images. To evaluate the effectiveness of this approach on inversion, we conducted tests where the generation process was conditioned by both historical frames and observed RTM image results. The inverted results are shown in Figure 7. We observed that the diffusion models can invert high-quality subsurface parameters, and the evolution of the inverted velocity, density, and CO2 saturation are reasonable and consistent. We also quantify the errors of the inverted results using the reference numerical simulations, shown in Figure 8. Compared to the reference results, our method showcases a good inversion performance in terms of accuracy, especially for the velocity and density models, which the images are more dependent on. This makes sense as the more conditions we have, like the RTM image results, which can provide strong constraints on velocity and density, the better the accuracy of the prediction of our method.
Another scenario of inversion is that we only have the information on permeability and observations like RTM images and try to estimate the CO2 saturation and elastic parameters. The inverted results using our proposed method are shown in Figure 9, and the quantitative analysis of the errors is shown in Figure 10. Besides good reconstruction of velocity and density changes, we find that the evolution of CO2 is better estimated as we have accurate permeability information. This validates the importance of the permeability to the inference of CO2 saturation.
Figure 9: Inversion results (b) given permeability and observed-data-based RTM images (a), which are denoted by the red box. The rows represent the various monitored properties described in Figure 4
@@ 3.5 Uncertainty analysis
Despite the generally high accuracy we managed to achieve in forecasting and inversion, we still had errors as compared to the numerical simulation the training of the diffusion models was based on. As we saw, the accuracy depends on the amount of data we use to condition the generation. Thus, uncertainty quantification as an additional output of the framework can help us assess confidence in the inversion and forecasting results in practical applications, where reference solutions are unavailable. Recall that our monitoring framework is based on the video diffusion model, which is one type of generative models that naturally supports the uncertainty analysis. The way to quantify approximate uncertainty within our framework is generally simple and cheap. We mainly need perform conditional sampling for the same history states and measurements starting from several random noise samples in parallel, let’s say 16 samples for the sake of the test though this number is theoretically small to properly sample the posterior distribution. Thus, we evaluate the mean and standard deviation of the generated 16 evolution samples. We assume a constant permeability in this test and give the three past history states and the RTM image results over the years to perform the inversion. We share one of the samples in Figure 12 as well as the mean prediction in Figure 13. Similar to the performance shown in Sections 3.2 and 3.3, the sample shows highly accurate inversion results, and the consistency of evolution between different variables remains. As for the mean values for those 16 samples, they are close to the reference numerical simulations as expected, but we found that the errors are slightly different compared to the sample shown in Figure 12b. They are larger at the area of the boundary of the CO2 plume and decrease from the boundary to the interior of the CO2 plume. This contributes to the uncertainty of the sampling process. The standard deviation values for those 16 samples are shown in Figure 14. As expected, high uncertainty appears at the boundary of the plume, especially around the leakage area, and decreases from the boundary to the interior of the CO2 plume, consistent with the errors map of the mean predictions. It demonstrates that the proposed method can achieve probabilistic predictions in a fully parallel way and can provide a reasonable approximate uncertainty map.
Figure 12: A sample (a) from the 16 inverted results given three history frames, constant permeability, and observed RTM images. (b) shows the errors between (a) and the reference numerical simulations.
Figure 13: The mean prediction of 16 inverted results (a) given three history frames, constant permeability, and observed RTM images starting from different random noises. (b) shows the errors between (a) and the reference numerical simulations.
@@
Figure 14: The standard deviation of 16 inverted results given three history frames, constant permeability, and observed RTM images starting from different random noises.
@@ 4 Dicussions
This paper presents a novel framework utilizing video diffusion models for subsurface forecasting and monitoring. It demonstrates good performance in terms of physical consistency, visual quality, and accuracy. However, a notable limitation is the computational cost of the sampling process, which necessitates iterative inference (1000 iterations in our case). This issue highlights the need for further optimization to reduce computational cost, possibly through approaches like denoising diffusion implicit models [Song et al., 2022] or consistency models [Song et al., 2023]. In addition, to handle the high-resolution multiphysics images over longer time lapse evolution, where the computational cost is expected to be huge, and the learning complexity increases, advanced techniques such as stable diffusion [Rombach et al., 2022] and [Peebles and Xie, 2022] are needed.
@@
Besides the computational cost, to apply our method in field applications, addressing the gap between the training data and realistic multiphysics evolution is required. Specifically, in realistic scenarios, we may have a time-dependent injection rate rather than a constant one, as used in this paper. In some cases, we might stop the injection for a while and continue later. What’s more, the location of the injection well and multiple injection wells should also be considered. Thus, including such a variable would be more interesting and important. The classifier-free conditional diffusion model [Ho and Salimans, 2021, Wang et al., 2024] would be a potential solution, which we might explore more in the future. For example, we can have the injection rate over a period as a condition, as such information is often known.
@@
In the current pipeline, we use RTM images as the seismic response rather than the seismic recordings. The adaption to seismic recordings also requires video diffusion models to learn a mapping from the velocity and density to the seismic wavefield. This mapping, as we know from full waveform inversion, is very complicated, requiring an abundance of training data, and often lacks generalization abilities. Although the ideal workflow is to enable the inversion or forecasting based on the observed seismic data, which is quite straightforward in realistic monitoring scenarios, such an undertaking will increase the complexity of the whole multiphysics process further as we need to teach the neural network to learn the wave-equation modeling engine. This challenge reveals another interesting future direction, in which we can try to learn a neural PDE solver [Huang et al., 2024, Huang and Alkhalifah, 2024b] for instant seismic modeling given the velocity and geometry or directly incorporate the differential seismic wavefield modeling engine, such as DeepWave [Richardson, 2023], into the monitoring framework.
@@
Unlike neural PDE solvers, which can learn simulation mappings between the given input multiple variables and the corresponding CO2 saturations, velocity, and so on, the proposed framework directly learns the end-to-end multiphysics evolution by processing them as video. Since the physical world and complex physical phenomena like CO2 multiphysics evolution can be hypothetically filmed and captured as a video, using a video diffusion model to store such evolution, including the physics involved, seems natural. Beyond the demonstration in this paper for this concept, there are also many literatures showing that the video diffusion models can simulate the world [Cachay et al., 2023, Price et al., 2023, Ding et al., 2024]. However, to make the diffusion models store the real physics and generate the evolution with arbitrary length, a large amount of training samples are needed. In addition, as shown in the paper, the forecasting results still exhibit errors compared to the reference results. To further enhance the performance and reduce the number of training samples, we believe that incorporating the physical knowledge, for example, the governing equations as a physics-constraint loss [Huang and Alkhalifah, 2022, 2024a, Huang et al., 2024, Christopher et al., 2024], in the sampling process can enhance the dynamic consistency over time, and correct the errors. To improve the forecasting ability to longer future horizons, combining the next-token prediction models, which enable variable-length generation, with the video diffusion models, e.g., diffusion forcing [Chen et al., 2024], might be a potential solution path for framework.
@@ 5 Conclusions
6 Acknowledgements

### v3

@@ 3.3 Inversion
In realistic monitoring scenarios, besides the historical information of the multiphysics evolution, we might also have some measurements of the seismic response on the surface, such as the seismic recordings, which can be converted to RTM images [Li and Li, 2021]. RTM images play a crucial role in these instances, acting as a strong conditioner that can constrain and guide the generation of CO2 evolution and the corresponding changes in subsurface elastic properties. With our method, we can do an inversion of the subsurface velocity, density, and CO2 saturation, given the observed seismic images. To evaluate the effectiveness of this approach on inversion, we conducted tests where the generation process was conditioned by both historical frames and observed RTM image results. The inverted results are shown in Figure 7. We observed that the diffusion models can invert high-quality subsurface parameters, and the evolution of the inverted velocity, density, and CO2 saturation are reasonable and consistent. We also quantify the errors of the inverted results using the reference numerical simulations, shown in Figure 8. Compared to the reference results, our method showcases a good inversion performance in terms of accuracy, especially for the velocity and density models, which the images are more dependent on. This makes sense as the more conditions we have, like the RTM image results, which can provide strong constraints on velocity and density, the better the accuracy of the prediction of our method. We also test the inversion performance on five distinct examples. The average relative L2 norm error is 0.021, and the average SSIM is 0.961.
Another scenario of inversion is that we only have the information on permeability and observations like RTM images and try to estimate the CO2 saturation and elastic parameters. The inverted results using our proposed method are shown in Figure 9, and the quantitative analysis of the errors is shown in Figure 10. They show good reconstruction of the velocity, density, and CO2 saturation evolutions. In such a scenario, the average relative L2 norm error of five distinct examples is 0.083, and the average SSIM is 0.984.
Figure 9: Inversion results (b) given permeability and observed-data-based RTM images (a), which are denoted by the red box. The solid color subfigures in (a) denote the unknown partial physical variables, which are the target of the inversion. The rows represent the various monitored properties described in Figure 4
@@ 3.5 Uncertainty analysis
Despite the generally high accuracy we managed to achieve in forecasting and inversion, we still had errors as compared to the numerical simulation the training of the diffusion models was based on. As we saw, the accuracy depends on the amount of data we use to condition the generation. Thus, uncertainty quantification as an additional output of the framework can help us assess confidence in the inversion and forecasting results in practical applications, where reference solutions are unavailable. Recall that our monitoring framework is based on the video diffusion model, which is one type of generative models that naturally supports the uncertainty analysis [Finzi et al., 2023, Price et al., 2023]. The way to quantify approximate uncertainty within our framework is generally simple and cheap. We mainly need perform conditional sampling for the same history states and measurements starting from several random noise samples in parallel, let’s say 16 samples for the sake of the test though this number is theoretically small to properly sample the posterior distribution. Thus, we evaluate the mean and standard deviation of the generated 16 evolution samples. We assume a constant permeability in this test and give the three past history states and the RTM image results over the years to perform the inversion. We show the reference numerical simulated results and the conditions used for generation in Figure 12. The mean prediction is shown in Figure 13a. Similar to the performance shown in Sections 3.2 and 3.3, the sample shows reasonable inversion results, and the consistency of evolution between different variables remains. The average relative error is 0.18, and the average SSIM is 0.9815. The standard deviation values for those 16 samples are shown in Figure 13b. As expected, high uncertainty appears at the boundary of the plume, especially around the potential leakage area, and decreases from the boundary to the interior of the CO2 plume. In addition, such uncertainty patterns might indicate regions where predictions are less reliable and thus require additional measurements, like well logs. Interpreting uncertainty in this context enables more informed decisions for taking preventive measures. Overall, the experiment demonstrates that the proposed method can achieve probabilistic predictions in a fully parallel way and can provide a reasonable approximate uncertainty map. Unlike ensemble approaches [Fort et al., 2020], which require training of multiple models, our approach can do the probabilistic generation within one model and is more efficient.
Figure 12: The reference numerical simulated results (a) and the given conditions used for the conditional generation (b). The solid color subfigures in (b) denote the unknown partial physical variables, which are the target of the inversion.
@@
Figure 13: The mean prediction of 16 inverted results (a) given three history frames, constant permeability, and observed RTM images starting from different random noises. (b) shows the standard deviation of these 16 inverted results.
@@ 4 Dicussions
This paper presents a novel framework utilizing video diffusion models for subsurface forecasting and monitoring. It demonstrates good performance in terms of physical consistency, visual quality, and accuracy. However, a notable limitation is the computational cost of the sampling process, which necessitates iterative inference (1000 iterations in our case). Generating one sample on a Nvidia A100 GPU will cost 29 seconds for unconditional generation and 84 seconds for conditional generation. This issue highlights the need for further optimization to reduce the computational cost, possibly through approaches like denoising diffusion implicit models [Song et al., 2022] or consistency models [Song et al., 2023]. In addition, to handle the high-resolution multiphysics images over longer time-lapse evolution, where the computational cost is expected to be huge, and the learning complexity increases, advanced techniques such as stable diffusion [Rombach et al., 2022] and [Peebles and Xie, 2022] are needed.
@@
Besides the computational cost, to apply our method in field applications, addressing the gap between the training data and realistic multiphysics evolution is required. Specifically, the permeability in this paper is assumed to be constant for simplification. Although incorporating this constant channel as additional information for different CO2 related multiphysics evolution has shown good performance in this paper, it might also have time-dependent properties in realistic scenarios. Hence, we can consider the time-dependent permeability in the training data preparation to make it better resemble a realistic scenario. The heterogeneous geological formations should also be considered during the data generation for complex field applications. Furthermore, we can add noise into the seismic recordings and adjust the seismic acquisition before obtaining the corresponding RTM images to imitate noisy real-world observations. In addition, we may have a time-dependent injection rate rather than a constant one, as used in this paper. In some cases, we might stop the injection for a while and continue later. What’s more, the location of the injection well and multiple injection wells should also be considered. Thus, including such a variable would be more interesting and important. The classifier-free conditional diffusion model [Ho and Salimans, 2021, Wang et al., 2024] would be a potential solution, which we might explore more in the future. For example, we can have the injection rate over a period as a condition, as such information is often known.
@@
In the current pipeline, we use RTM images as the seismic response rather than the seismic recordings. The adaption to seismic recordings also requires video diffusion models to learn a mapping from the velocity and density to the seismic wavefield. This mapping, as we know from full waveform inversion, is very complicated, requiring an abundance of training data, and often lacks generalization abilities. Although the ideal workflow is to enable the inversion or forecasting based on the observed seismic data, which is quite straightforward in realistic monitoring scenarios, such an undertaking will increase the complexity of the whole multiphysics process further as we need to teach the neural network to learn the wave-equation modeling engine. This challenge reveals another interesting future direction, in which we can try to learn a neural PDE solver [Huang et al., 2024, Huang and Alkhalifah, 2024] for instant seismic modeling given the velocity and geometry or directly incorporate the differential seismic wavefield modeling engine, such as DeepWave [Richardson, 2023], into the monitoring framework.
@@
Unlike neural PDE solvers, which can learn simulation mappings between the given input multiple variables and the corresponding CO2 saturations, velocity, and so on, the proposed framework directly learns the end-to-end multiphysics evolution by processing them as video. Since the physical world and complex physical phenomena like CO2 multiphysics evolution can be hypothetically filmed and captured as a video, using a video diffusion model to store such evolution, including the physics involved, seems natural. Beyond the demonstration in this paper for this concept, there are also many literatures showing that the video diffusion models can simulate the world [Cachay et al., 2023, Price et al., 2023, Ding et al., 2024]. However, to make the diffusion models store the real physics and generate the evolution with arbitrary length, a large amount of training samples are needed. In addition, as shown in the paper, the forecasting results still exhibit errors compared to the reference results. To further enhance the performance and reduce the number of training samples, we believe that incorporating the physical knowledge, for example, the governing equations as a physics-constraint loss [Huang et al., 2024, Christopher et al., 2024], in the sampling process can enhance the dynamic consistency over time, and correct the errors. To improve the forecasting ability to longer future horizons, combining the next-token prediction models, which enable variable-length generation, with the video diffusion models, e.g., diffusion forcing [Chen et al., 2024a], might be a potential solution path for the framework.
@@ 5 Conclusions
6 Open Research Section
The datasets used in this paper can be downloaded at Huang [2025]. During the dataset generation, we use Jutul.jl Møyner et al. [2023] and JutulDarcyRules.jl Yin et al. [2023a] for flow simulation and JUDI Louboutin et al. [2023a] for RTM images generation. Codes of the video diffusion models to reproduce the results presented here are also available at Huang [2025]. Figures were made with Matplotlib version 3.2.1 Caswell et al. [2020].
7 Acknowledgments